Al calculated the profit from his candy shop is given by the formula P = - 10x2+ 100x + 300, where x is the number of units of candies sold daily. How many units should be sold for a profit of $460?
Answer by ikleyn(52781)

.

To answer this question, you need to solve quadratic equation


    -10x^2 + 100x + 300 = 460.


It is the same as


    10x^2 - 100x + 160 = 0,  or


      x^2 - 10x + 16 = 0


Factor left side


      (x-8)*(x-2) = 0.


The roots are 8 and 2; they both are good as the solutions.



ANSWER.  There are 2 possible answers: 8 units of candies and/or 2 units of candies.



CHECK.  You may (and you should/must) check both answers by substituting the values into the given original formula.

        I did it for myself (!)   You do it for yourself (!)

Solved.